# Quantitative Comparison of Fine-Tuning Techniques for Pretrained Latent Diffusion Models in the Generation of Unseen SAR Image Concepts

### v1

@@ 5.3 Study on LoRA Rank and Alpha
Figure 8: Mean cosine distance between image-text pairs, compared to real ones.
Contrary to what has been observed in other domains, our results suggest that increasing the LoRA rank and alpha beyond moderate values does not improve vision–language alignment in the SAR generation setting. As shown in Table 4, the best CLIP alignment metrics (lowest rank mean and variance) are achieved with relatively low-rank settings (r = 8, alpha = 4), as used in the soleil-up-7 model. Larger ranks, such as those in fiber-network-6 and screen-light-4, lead to slightly degraded alignment scores despite similar KL divergence.
Figure 8 further confirms this: the cosine similarities between image–prompt pairs for soleil-up-7 are closer to those observed between real SAR images and their captions, indicating more faithful semantic representation. In contrast, higher LoRA configurations tend to blur category separability, possibly due to overfitting or excessive parameter capacity in the text encoders.
These results suggest that compact LoRA adapters (with low rank and scaling) are sufficient for prompt-image alignment in this domain and may generalize better than overparameterized configurations.
@@ 5.4 Study on Batch Size
Table 5: Comparison of batch size at epoch 8 with a fixed training configuration (UNet full fine-tuning F, LoRA for Text Encoders).
@@
As shown in Table 5 and Figure 9, increasing the batch size generally improves the similarity between generated and real SAR distributions, as measured by the KL divergence. Larger batches expose the model to a broader diversity of textures and amplitude statistics, helping it learn more robust and representative features. The best KL score is observed for the 64-image configuration, suggesting a potential trade-off: extremely large batches (e.g., 128) may lead to diminishing returns or underfitting, while mid-sized batches offer an optimal balance between generalization and convergence.
@@ 5.5 Study on VAE Decoder
While most studies using SDXL retain the original Variational Autoencoder (VAE) without modification, we hypothesize that fine-tuning its encoder and/or decoder components may improve latent representation quality for SAR images. This is motivated by the fact that SAR data differ significantly from natural images in terms of statistical structure, noise characteristics (e.g., speckle), and semantic content.
Fine-tuning the SDXL VAE, however, proves particularly challenging when transferring to a new domain such as SAR. In our experiments, fine-tuning the VAE encoder often led to latent space instability and mode collapse. In practice, the pretrained VAE already reconstructs SAR images with low error, and the decoded outputs follow a Rayleigh-like amplitude distribution that closely resembles real SAR data. Attempts to adjust the encoder disrupted this alignment and led to overfitting.
Moreover, training separately of the VAE from the rest of the model (UNet and text encoders) proved to be ineffective, because coherence across components is essential in foundation model pipelines. To address this, we chose to fine-tune only the VAE decoder jointly with the UNet and both text encoders during a final refinement phase.
Despite the VAE’s strong reconstruction ability in pixel space, generating SAR images from pure Gaussian noise during inference remains challenging. While UNet fine-tuning helps, the decoder often fails to fully reproduce the sharp textures and amplitude dynamics typical of SAR data.
To mitigate this, we performed a short refinement training of the decoder (along with the UNet and text encoders)
$z_{0}=\frac{1}{\sqrt{\bar{\alpha}_{t}}}\left(\hat{z}_{t}-\sqrt{1-\bar{\alpha}_{t}}\cdot\epsilon\right)$ (9)
where $\bar{\alpha}_{t}$ is the cumulative product of noise schedule coefficients up to timestep $t$. The resulting $z_{0}$ is decoded into a synthetic SAR amplitude image $\hat{x}$ via the VAE decoder.
To ensure that $\hat{x}$ approximates real SAR distributions, we include a KL divergence term comparing the pixel amplitude histograms of $\hat{x}$ and the ground-truth image $x$:
$\mathcal{L}_{\text{KL}}=D_{\text{KL}}(P_{\text{real}}(x)\parallel P_{\text{gen}}(\hat{x}))$ (10)
The total loss used for decoder fine-tuning is:
$\mathcal{L}_{\text{total}}=\mathcal{L}_{\text{Rayleigh}}+\mathcal{L}_{\text{KL}}$
This formulation encourages the decoder to better approximate SAR-specific amplitude statistics during generation.
@@
Figure 10: Comparison of KL distance distributions between 330 real and generated flattened images - with 833 refining training steps - with 40 timesteps with 15 timesteps on the last 15 % of the denoising process.
@@
Figure 11: Comparison of generated images — at epoch 8 — with the same prompts: (1) "A satellite view of a port with a boat in the water and a forest nearby." (2) "A satellite view of a vast expanse of land with a circular structure and a few isolated buildings." (3) "A satellite view of a dense forest with a river in the center of the forest."
Overall, we observe that the model whale-north-8-refined yields better semantic coherence and finer structural details. As illustrated in Figure 11, it successfully adds a boat to the harbor, generate a circular structure in an open field, and delineates a river through the center of a forest — all in accordance with the textual prompts. The decoder was refined over 833 steps (equivalent to one epoch), limited intentionally to avoid latent space collapse, which had previously manifested as color artifacts in the generated SAR images. Thus, training the VAE remains delicate, as even minor instabilities can propagate to the generation process and degrade structural or statistical fidelity.
@@ 5.6 Study on <SAR> Token Embedding Learning
To improve the model’s transferability to the SAR domain, we adopt a token embedding learning strategy inspired by textual inversion. Instead of relying solely on full model fine-tuning, we introduce a new learnable token <SAR> whose embedding is optimized jointly with the model weights.
During training, we modify the prompts by replacing general descriptions like “A satellite view of…” with SAR-specific expressions such as “A <SAR> image of…”. This encourages the model to associate the <SAR> token with the statistical and structural patterns specific to radar imagery. The embedding is updated through backpropagation along with the UNet, aligning the textual representation with SAR-specific latent visual features.
@@
The model heart-rose-2 achieves the lowest KL divergence (0.23) and optimal CLIP Rank scores (1.68 and 1.31 for $\nu$ and $\sigma$, respectively), indicating superior performance in generating realistic SAR images (Table 7). Histogram comparison (Figure 12) reveals that the probability density distribution of heart-rose-2 more closely captures the inherent dynamics of SAR data, indicating that this model learns better representational dynamics of Synthetic Aperture Radar imagery.
@@
Table 8: Per-category CLIP Rank statistics ($\nu$, $\sigma$) and KL divergence for the models whale-north-8 and heart-rose-2.
@@
Figure 13: Comparison of generated images — at epoch 8 — with the same prompts: (1) "A SAR image of a dense forest with a river in the center of the forest." (2) "A SAR image of a vast expanse of land with a circular structure and a few isolated buildings."
As shown in Table 8, the heart-rose-2 model outperforms whale-north-8 across nearly all categories, especially in forested, mountainous, and seacoast scenes — which typically exhibit more complex scattering patterns. Notably, it is capable of composing scenes that were never explicitly observed during training like “a circular structure in a city center” or “a river in the middle of a dense forest” (Figure 13). In general, heart-rose-2 generates high-quality SAR images with realistic spatial structures and texture patterns, as illustrated in Figure 20 (see Appendix A.3).
@@
As detailed in Section 4.3, GLCM-based texture metrics provide further evidence of generative realism. Figure 14 shows that the model effectively reproduces directional patterns for contrast, entropy, and homogeneity. However, deviations remain in the correlation metric, especially at certain angles, suggesting that the model may not fully capture the spatial dependencies induced by SAR imaging geometry or coherent scattering mechanisms.
@@ 6 Applications
Our fine-tuned Stable Diffusion XL model for SAR imagery can be used in several practical applications. One of its key strengths lies in its ability to perform image-to-image generation while preserving essential statistical properties of SAR data, such as speckle texture and reflectivity distributions. This makes it particularly useful for denoising, adding spatial detail, or enhancing outputs from ONERA’s physics-based simulators.
In previous work, we demonstrated the effectiveness of our early fine-tuned model in a conditional, multi-resolution ControlNet pipeline for large-scale scene generation [29]. We also showed its ability to transform satellite TerraSAR-X data into high-resolution, airborne-like imagery with reduced sensor noise [30]. These applications rely on ControlNet to guide generation using structural priors.
Here, we illustrate two new use cases: enhancing real SAR images and improving synthetic outputs from physics-based simulators.
6.1 Enhancing TerraSAR-X image at 40cm
@@
6.2 Enhancing physics-based simulator image: subsampled from 80cm to 40cm
We also apply our model to improve the realism of synthetic SAR images generated by ONERA’s EMPRISE simulator. These images are enhanced to include finer textures and more realistic spatial structure.
@@ 7 Discussion and Conclusion
This work is, to our knowledge, the first comparison study of fine-tuning techniques on a large-scale latent diffusion model — Stable Diffusion XL (SDXL) for a radically different imaging modality: Synthetic Aperture Radar (SAR). While most existing research focuses on generative adaptation for natural images (optical, RGB, photorealistic domains), we show that it is possible to adapt a vision-language foundation model to generate physically grounded, speckle-heavy SAR data. This contribution goes beyond SAR itself: it opens up new possibilities for adapting foundation models to other underrepresented sensing domains such as hyperspectral, infrared, depth, or medical imaging.
By exploring multiple fine-tuning strategies—full UNet updates, Low-Rank Adaptation (LoRA) for text encoders, token embedding learning, and VAE decoder refinement, we provide a reproducible framework to adapt large pretrained diffusion models to unseen imaging physics. We further propose evaluation tools specific to SAR, including statistical distance metrics (KL divergence), structural indicators (GLCM features), and a vision-language alignment score based on a CLIP model fine-tuned for SAR.
However, adapting SDXL to SAR comes with key challenges. First, the statistical and visual properties of SAR data are fundamentally different from natural images: they are grayscale, noisy (speckle), anisotropic, and heavily dependent on acquisition geometry. Foundation models pretrained on optical data tend to favor smooth textures, uniform lighting, and naturalistic shapes — biases that can degrade realism in generated SAR samples.
Second, our dataset, although sizeable (100K image-caption pairs), remains imbalanced. Some scene types (e.g., forest or water) dominate the distribution, while others (e.g., airport or industrial zones) are underrepresented, making fine-tuning harder. Furthermore, since SAR lacks direct semantic interpretation, we rely on captions generated from optical counterparts using CogVLM2. While this provides rich textual supervision, it introduces hallucinations due to mismatched spatial coordinates between optical and SAR images.
Normalization is another open issue. We scaled SAR amplitudes to [0,1] using a clipping strategy that introduced an artificial saturation peak, which complicates the generative process. This calls for more principled pre-processing methods that preserve SAR-specific statistics, such as using phase and frequency-domain data, or learning from complex-valued representations directly.
Evaluation of SAR generation remains a bottleneck. Conventional metrics like FID and CLIP score are inadequate for SAR due to their optical training biases. Although we fine-tuned a CLIP model on SAR-caption data, assessing realism and structure fidelity is still difficult without standardized benchmarks or human-perceptual baselines tailored to SAR data.
Despite these limitations, our experiments demonstrate that:
  • 1.
    Full UNet fine-tuning outperforms LoRA-only approaches in capturing SAR-specific structures.
  • 2.
    Embedding a learned token like <SAR> improves semantic alignment and generation quality.
  • 3.
    Fine-tuning the VAE decoder — limited to later timesteps in the reverse diffusion — enhances textural realism without destabilizing the latent space.
  • 4.
    Generated samples can successfully enhance real satellite or simulated SAR images, suggesting practical use for denoising, simulation refinement, and data augmentation.
@@
In summary, this work sets a precedent for adapting pretrained generative models to unconventional data domains. It delivers insights for both the SAR and broader AI communities on how to transfer powerful foundation models to non-optical, physics-driven settings. Future directions include: multimodal conditioning (e.g., combining text with elevation or optical inputs), learning in the complex SAR domain (amplitude + phase), and developing SAR - native perceptual metrics to better capture realism, structure, and semantic consistency.
Appendix A Appendix
A.1 Mean Absolute Weight Change (MAWC)
@@
$\Delta w_{i}=\left|w_{i}^{(1)}-w_{i}^{(0)}\right|$ (11)
@@
$\mathrm{MAWC}=\frac{1}{W}\sum_{i=1}^{W}\Delta w_{i}$ (12)
@@
Figure 17: Mean Absolute Weight Difference per block and sub-block of the UNet during the training configuration soleil-up-7 with a resolution threshold of $5\times 10^{-4}$
A.2 Training images generation
To visually assess the quality and diversity of generations, we display side-by-side image samples produced by each model across six representative scene categories.

### v2

@@ 5.3 Effect of LoRA Rank and Scaling
Figure 9: Mean cosine distance between image-text pairs, compared to real ones using SAR-CLIP Model.
Contrary to what has been observed in the literature, our results suggest that increasing the LoRA rank beyond certain values does not enhance vision–language alignment in the SAR generation setting. As shown in Table 3, the best (lowest) CLIP Rank score is achieved with a relatively low-rank setting ($r=8$, $\alpha=4$) used in the soleil-up-7 model, whereas larger ranks (e.g., fiber-network-6, screen-light-4) do not produce higher scores.
From the equation below, and since we set $\alpha=r/2$, increasing $r$ increases the adapter parameter count $r(m+n)$ and raises the capacity of $\Delta W$ (because $\mathrm{rank}(\Delta W)\leq r$), while the effective update scale $\alpha/r$ remains constant. The results indicate that, for text-encoder adaptation to SAR, adding degrees of freedom to $\Delta W$ beyond a moderate level does not translate into better vision–language alignment on our data; the low-rank setting preserves alignment more effectively for the same update magnitude.
Figure 9 is consistent with this interpretation: cosine similarities between image–prompt pairs for soleil-up-7 are closer to those for real SAR–caption pairs, indicating more faithful semantic alignment under low-rank updates.
@@ 5.4 Study on Batch Size
Table 4: Comparison of batch size values at epoch 8 with a fixed training configuration (UNet full fine-tuning F, LoRA for Text Encoders).
@@
As shown in Table 4 and Figure 10, increasing the batch size generally improves the similarity between generated and real SAR distributions, as measured by the KL divergence. Larger batches expose the model to a broader diversity of textures and amplitude statistics, helping it learn more robust and representative features. The best KL score is observed for the whale-north-8 model (batch size = 64), suggesting a potential trade-off: extremely large batches (e.g., 128) may lead to diminishing returns or underfitting, while mid-sized batches offer an optimal balance between generalization and convergence.
@@ 5.5 Study on VAE Decoder
While most studies using SDXL take the original Variational Autoencoder (VAE) without modification, we hypothesize that fine-tuning its encoder and/or decoder components may improve latent representation quality for SAR images. This is motivated by the fact that SAR data differ significantly from natural images in terms of statistical structure, noise characteristics (e.g., speckle), and semantic content.
Fine-tuning the pretrained VAE used in SDXL, however, is particularly challenging when transferring to a new domain such as SAR. In our experiments, fine-tuning the VAE encoder often led to latent space instability and mode collapse. In practice, the pretrained VAE already reconstructs SAR images with low error, and the decoded outputs follow a Rayleigh-like amplitude distribution that closely resembles real SAR data. Attempts to adjust the encoder disrupted this alignment and led to overfitting.
Moreover, training separately the VAE from the rest of the model (UNet and text encoders) proved to be ineffective, because coherence across components is essential in foundation model pipelines. To address this, we chose to fine-tune only the VAE decoder jointly with the UNet and both text encoders during a final refinement phase.
Despite the VAE’s strong reconstruction ability in pixel space, generating SAR images from pure Gaussian noise during inference remains challenging. While UNet fine-tuning helps, the decoder often fails to fully reproduce the textures and amplitude dynamics typical of SAR data.
To mitigate this, we perform a short refinement of the VAE decoder together with the UNet and the text encoders. As noted in Section 4.1, the model is conditioned with timesteps $t\in[0,1000]$ sampled uniformly. To improve performance while preserving the pretrained semantic prior, we bias fine-tuning toward the final stages of the reverse diffusion (low-noise regime), which correspond to early timesteps of the forward process. Concretely, we run a single epoch and restrict training timesteps to the last 15% of the reverse-diffusion schedule, which concentrates learning near reconstruction and reduces latent drift, improving SAR image fidelity.
In addition, we add a Kullback–Leibler (KL) divergence term to the loss to minimize the divergence between the amplitude distribution of the generated image $\hat{x}$ and that of the target image $x$:
$).\mathcal{L}_{\mathrm{KL}}\;=\;D_{\mathrm{KL}}\!\big{(}P_{\mathrm{real}}(x)\,\|\,P_{\mathrm{gen}}(\hat{x}$$\big{)}$. (10)
The total loss used during refinement is
$\mathcal{L}_{\mathrm{refine}}\;=\;\mathcal{L}_{\mathrm{base}}\;+\;\lambda_{\mathrm{KL}}\,\mathcal{L}_{\mathrm{KL}},$ (11)
where $\mathcal{L}_{\mathrm{base}}$ denotes the standard diffusion (noise-prediction) loss used in our setup, and $\lambda_{\mathrm{KL}}$ controls the weight of the distribution-matching term.
@@
Figure 11: Comparison of KL distance distributions between 330 real and generated flattened images - with a refining training strategy on the last 15 % of the denoising process.
@@
Figure 12: Comparison of generated images (1024x1024px at 40cm) — at epoch 8 — with the same prompts: (1) "A satellite view of a port with a boat in the water and a forest nearby." (2) "A satellite view of a vast expanse of land with a circular structure and a few isolated buildings." (3) "A satellite view of a dense forest with a river in the center of the forest."
Overall, we observe that the model whale-north-8-refined has both SAR-specific fidelity and prompt-conditioned semantic competence. Beyond producing SAR imagery, it uses semantic knowledge of the pretrained base model to assemble out-of-distribution scene configurations. As illustrated in  12, it successfully adds a boat within a harbor, synthesizes a circular installation in open terrain, and delineates a river through a forest—each consistent with the textual prompt in a manner that is physically consistent with SAR imaging.
@@ 5.6 Study on <SAR> Token Embedding Learning
To improve transfer to the SAR domain, we adopt a token-learning strategy inspired by textual inversion. We extend the tokenizer vocabulary with a new token <SAR> and assign it a learnable embedding optimized jointly with the model parameters.
During training, we modify the prompts by replacing general descriptions like “A satellite view of…” with SAR-specific expressions such as “A <SAR> image of…”. This encourages the model to associate the <SAR> token with the statistical and structural patterns specific to radar imagery. Gradients from the diffusion objective (and our distributional terms) update the token embedding together with the UNet and the text encoders, aligning the textual representation with SAR-consistent latent visual features.
@@
The model heart-rose-2 achieves the lowest KL divergence (0.23) and optimal CLIP Rank scores (1.68 and 1.31 for $\nu$ and $\sigma$, respectively), indicating superior performance in generating realistic SAR images (Table 6). Histogram comparison (Figure 13) reveals that the probability density distribution of heart-rose-2 more closely captures the inherent dynamics of SAR data, indicating that this model learns better representational dynamics of Synthetic Aperture Radar imagery. And, it is capable of composing scenes that were never explicitly observed during training like “a circular structure in a city center” or “a river in the middle of a dense forest” (Figure 14). In general, heart-rose-2 generates high-quality SAR images with realistic spatial structures and texture patterns, as illustrated in Figure 16.
@@
Table 7: Per-category CLIP Rank scores and KL divergence for two models whale-north-8 and heart-rose-2.
As shown in Table 7, the heart-rose-2 model outperforms whale-north-8 in nearly all categories, particularly in forested, mountainous, and seacoast scenes, which typically exhibit more complex scattering patterns. Specifically, the heart-rose-2 model achieves the lowest KL distance across all categories and the best CLIP Rank scores in 8 out of 11 categories.
@@
Figure 14: Comparison of generated images (1024x1024px at 40cm) — at epoch 8 — with the same prompts: (1) "A SAR image of a dense forest with a river in the center of the forest." (2) "A SAR image of a vast expanse of land with a circular structure and a few isolated buildings."
As detailed in Section 4.3, GLCM-based texture metrics provide further insight into the model’s generative realism. Figure 15 shows that directional patterns, particularly for metrics like entropy, are well reproduced. However, contrast is less accurate across all angles. Indeed, in Figure 13, the KL divergence of the heart-rose-2 model shows that the generated pixel distribution is smoother compared to the real Rayleigh distribution, with fewer black pixels (low-intensity) and more white pixels (high-intensity) in the generated images. In terms of spatial correlation of texture, which corresponds to large-scale spatial dependencies related to SAR geometry, the model heart-rose-2 performs well, although some variations are observed across different angles.
@@ 6 Applications
Our fine-tuned Stable Diffusion XL model for SAR imagery offers several practical applications. A significant advantage of this model is its ability to generate novel data beyond the training domain, creating new and unique content not present in the original dataset.
Figure 16: Generated images (1024x1024px at 40cm) from the heart-rose-2 model. Each image corresponds to a distinct textual prompt (see Appendix C).
As illustrated in Figure 16, we observe that we are able to generate various types of landscapes, such as fields, forests, seacoasts, or mountains, that the model has never seen during its training. This includes scenarios like a boat near a forested coastline (c) or a bridge with a particular design in the middle of a landscape (k). These results demonstrate the model’s capacity to produce creative and realistic scenarios from its language understanding.
This is possible because the model, based on a pretrained architecture, is capable of producing variations of objects or rare situations that are not overfitted to the limited examples seen during training. Consequently, the model can generate realistic and representative variants of these uncommon cases or objects, making it especially effective for tasks requiring creative generation of new SAR imagery.
For example, it can be used to perform image-to-image generation while preserving essential statistical properties of SAR data, such as speckle texture and reflectivity distributions. This makes the model particularly valuable for applications like adding spatial detail, or refining outputs from ONERA’s physics-based simulators.
More details can be found in our previous work (Debuysère et al. (2024), Debuysère et al. (2025) and Trouve et al. (2024)) where we demonstrated the effectiveness of our fine-tuned model in a conditional, multi-resolution ControlNet pipeline for large-scale scene generation. We also showed its ability to transform satellite TerraSAR-X data into high-resolution, airborne-like imagery with reduced sensor noise. These applications rely on ControlNet to guide generation using structural priors.
In this section, we illustrate two use cases: TerraSAR-X conditioned 40 cm synthesis and simulator-conditioned 80 cm synthesis.
6.1 TerraSAR-X–Conditioned Synthesis at 40 cm
@@
6.2 Simulated SAR Image-Conditioned Synthesis
We also apply our model to ONERA’s EMPRISE outputs (80 cm). The pipeline produces 40 cm imagery with finer textures and more realistic spatial structure.
@@ 7 Discussion and Conclusion
We presented, to our knowledge, the first comparative study of fine-tuning strategies for a large latent diffusion model—Stable Diffusion XL (SDXL)—on Synthetic Aperture Radar (SAR) imagery. While most prior work addresses optical domains, our results show that a vision–language foundation model can be adapted to generate physically grounded SAR scenes.
Experiments indicate that full UNet fine-tuning is most effective for learning SAR-specific structure, whereas text-encoder adapters benefit from low-rank updates and a learned <SAR> token to preserve prompt fidelity. A brief low-noise refinement of the VAE decoder further improves textural realism without destabilizing the latent space. Beyond unconditional synthesis, the model supports conditioned generation for practical use cases, including TerraSAR-X–guided 40 cm synthesis and refinement of physics-based simulator outputs.
This capability enables scalable data augmentation and composition of rare or operationally relevant scenarios. Leveraging open-source foundation models also facilitates multimodal conditioning (e.g., segmentation, depth) and ControlNet-based guidance.
Limitations include evaluation restricted to X-band at 40cm slant-range sampling distance and specific processing settings. Nonetheless, although our experiments target 40cm, we detail a parameter-efficient procedure to adapt the model to other resolutions using low computational resources (one GPU H100). Future directions include multimodal conditioning with elevation/optical inputs and extending learning to the complex domain (amplitude and phase).
@@
In summary, this work highlights the potential of using generative AI for large-scale SAR scene generation, a domain where physics-based simulators are still underdeveloped. It sets a precedent for adapting pretrained generative models to unconventional data domains. It delivers insights for both the SAR and broader AI communities on how to transfer powerful foundation models to non-optical, physics-driven settings. Future directions include: multimodal conditioning (e.g., combining text with elevation or optical inputs), and learning in the complex SAR domain (amplitude and phase).
8 CRediT authorship contribution statement
Declaration of generative AI and AI-assisted technologies in the writing process
During the preparation of this work the authors used ChatGPT to assist with the translation from French to English. After using this tool, the authors reviewed and edited the content as needed and take full responsibility for the content of the publication.
Declaration of competing interest
The authors declare that they have no known competing financial interests or personal relationships that could have appeared to influence the work reported in this paper.
9 Acknowledgments
This work was supported by ONERA - The French Aerospace Lab, with financial support from the French Ministry of Defence (MoD). This research was conducted as part of Solène Debuysère’s PhD thesis. The authors wish to express their sincere gratitude to all those who contributed to the success of this work.
Appendix A Mean Absolute Weight Change (MAWC)
@@
$\Delta w_{i}=\left|w_{i}^{(1)}-w_{i}^{(0)}\right|$ (A.1)
@@
$\mathrm{MAWC}=\frac{1}{W}\sum_{i=1}^{W}\Delta w_{i}$ (A.2)
@@
Figure A.1: Mean Absolute Weight Difference per block and sub-block of the UNet during the training configuration soleil-up-7 with a resolution threshold of $5\times 10^{-4}$
Appendix B Training images generation
To visually assess the quality and diversity of generations, we display side-by-side image samples produced by each model across six representative scene categories at epoch 8 using the same seed.